### v1

# Effective (Floquet) Lindblad generators from spectral unwinding

Abstract
A mathematical description of the reduced dynamics of an open quantum system can often be given in terms of a completely positive and trace preserving (CPTP) map, also known as quantum channel. In a seminal work by Wolf et al. [Phys. Rev. Lett. 101, 150402 (2008)], it was shown that deciding whether a given quantum channel was generated from an underlying effective Markovian dynamics, with time-independent Lindbladian generator, is generally an NP-hard problem. In this work we show that in cases where one has access to the full reduced dynamics at all previous times (the dynamical map) one can significantly facilitate the search for an effective generator by making use of Floquet theory. By performing a spectral unwinding such that the effective micromotion is minimized, the effective Floquet generator is often an excellent candidate for an effective Lindblad generator, hence significantly reducing the complexity of the search for an effective Lindblad generator in many (though not all) cases. Our results are relevant for engineering Floquet Lindbladians in complex many-body systems.
Introduction.—Digital and analog simulation of quantum many-body systems have yielded fascinating results [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Both rely in part on the advent of Floquet engineering [19, 20, 5, 6, 21] (which in the context of digital simulation is usually referred to as trotterization [22, 23, 24, 25, 26, 27, 28, 29]). The central idea is that for an isolated quantum system with time-periodic Hamiltonian $H(t)=H(t+T)$, the one-cycle time-evolution operator $U(T)$ can be rewritten as $U(T)=\exp(-iH_{\mathrm{F}}T)$ with a time-independent Floquet Hamiltonian $H_{\mathrm{F}}$ with possibly novel terms that are not present in the lab Hamiltonian, e.g. artificial gauge fields [3, 4, 30, 19, 5, 6, 31, 32, 33] or nontrivial interactions [34, 35, 36, 37, 13, 38, 39].
Recently, the question was raised [40, 41, 42] whether these ideas can be generalized to open Floquet systems [43], i.e. that interact with an environment (e.g. by coupling the system to external photon or particle reservoirs). Interestingly, it was shown [41] that the dynamics under a time-periodic Lindbladian $\mathcal{L}(t)=\mathcal{L}(t+T)$ [41, 27, 44, 45, 46, 47] cannot always be understood as generated from an effective time-independent Floquet Lindbladian $\mathcal{L}_{\mathrm{F}}$ [40, 48, 49, 41, 42, 50, 51, 44, 52]. This is due to effective non-Markovian effects that are built up during the evolution. Nevertheless, as a result of the existence of branches when calculating the complex logarithm of the dynamical map, deciding whether a valid Floquet Lindbladian exists is generally an NP-hard problem [53]. This question closely relates to earlier work by Wolf et al. [54, 53] where the existence of an effective Lindbladian generator was studied for a general completely positive and trace-preserving (CPTP) map $\mathcal{V}$. In this work we show that, by using ideas from Floquet theory, the general problem of deciding Markovianity can be addressed more efficiently if the full dynamical map $\mathcal{V}(t)$ is known for all times $t$. Our results lay the foundation for Floquet engineering of Lindbladians in extended many-body systems.
Quantum Channels, Dynamical maps.—The mathematical cornerstone of the theory of Markovian open quantum systems are one-parameter semigroups $\mathcal{V}(t)\in L(L(\mathcal{H}))$, where $L(\mathcal{H})$ denotes the space of linear operators on the Hilbert space $\mathcal{H}$ with dimension $d$. The semigroup $\mathcal{V}(t)=e^{\mathcal{L}t}$, $t\geq 0$, describes an evolution of the density operator $\varrho(t)=\mathcal{V}(t)[\varrho(0)]$ that is Markovian and consistent with a physical evolution, if the dynamical map $\mathcal{V}(t)$ is CPTP, i.e. a quantum channel, at all times $t$. Lindblad [55], Gorini, Kossakowski and Sudarshan [56] have shown that a superoperator $\mathcal{L}$ is the generator of a quantum dynamical semigroup $\mathcal{V}(t)=e^{\mathcal{L}t}$, iff it can be expressed as
$\mathcal{L}[\cdot]=-{i}\Bigl{[}H,\cdot\Bigr{]}+\sum_{i,j=1}^{d^{2}-1}G_{ij}\biggl{(}F_{i}\cdot F_{j}^{\dagger}-\frac{1}{2}\Bigl{\{}F_{j}^{\dagger}F_{i},\cdot\Bigr{\}}\biggr{)},$ (1)
[…]
Markovianity test.—Let us turn to the converse question and ask “Given an arbitrary quantum channel $\mathcal{V}$ (say at time $T$) is there an effective generator $\mathcal{S}=\log(\mathcal{V})/T$ that has Lindblad form as in Eq. (1)?”. Due to the multi-valuedness of the complex logarithm, this question turns out to result in an NP-hard problem in general, also called the Markovianity problem [54, 53]. An algorithm to address this problem was proposed by Wolf et al. [54, 53]: A generator $\mathcal{S}\in L(L(\mathcal{H}))$ has Lindblad form, iff (i) for all Hermitian $\kappa=\kappa^{\dagger}$, $\mathcal{S}$ preserves Hermiticity, i.e. $\mathcal{S}[\kappa]=\bigl{(}\mathcal{S}[\kappa]\bigr{)}^{\dagger}$, and (ii) $\mathcal{S}$ is conditionally completely positive (CCP), i.e.
$(\mathbf{1}_{d^{2}}-\Sigma)\mathcal{S}^{\Gamma}(\mathbf{1}_{d^{2}}-\Sigma)\geq 0.$ (2)
Here, $\Sigma={|{\Omega}\rangle}{\langle{\Omega}|}$ is the projector onto the maximally entangled state ${|{\Omega}\rangle}=\frac{1}{\sqrt{d}}\sum_{j=1}^{d}{|{j}\rangle}\otimes{|{j}\rangle}$ of the system and an ancilla of same dimension $d$. $\mathcal{S}^{\Gamma}=d\cdot(\mathcal{S}\hskip 1.42262pt\otimes\hskip 1.42262pt\mathbf{1}_{d})[\Sigma]$ denotes the Choi representation of $\mathcal{S}$. Since the complex logarithm is not uniquely defined, we need to check if there is a branch of the operator logarithm of $\mathcal{V}$ that satisfies condition (i) and (ii).
To do this, we introduce the scalar product $\langle\!\langle A|B\rangle\!\rangle=\operatorname{Tr}[A^{\dagger}B]$ on the Hilbert space $L(\mathcal{H})$. By also defining an orthonormal basis $\{F_{\alpha}\}_{\alpha=1}^{d^{2}}$ in $L(\mathcal{H})$, we can represent quantum channels $\mathcal{V}$ as matrices $\hat{\mathcal{V}}$, with their corresponding matrix elements $\hat{\mathcal{V}}_{\alpha\beta}={\langle\!\langle{F_{\alpha}}|}\mathcal{V}{|{F_{\beta}}\rangle\!\rangle}$. On this space, (density) matrices $\varrho={|{i}\rangle}{\langle{j}|}$ become vectors ${|{\varrho}\rangle\!\rangle}={|{i,j}\rangle\!\rangle}$, hence this identification is known as vectorization of the density matrix. We first cast the map $\hat{\mathcal{V}}$ into Jordan normal form [54], $\hat{\mathcal{V}}=\sum_{r}\lambda_{r}{|{r_{r}}\rangle\!\rangle}{\langle\!\langle{l_{r}}|}+\sum_{c=1}^{N_{c}}\bigl{(}\lambda_{c}{|{r_{c}}\rangle\!\rangle}{\langle\!\langle{l_{c}}|}+\lambda_{c}^{*}\ \mathbb{F}{|{r_{c}}\rangle\!\rangle}{\langle\!\langle{l_{c}}|}\mathbb{F}\bigr{)}$, with $r$ indexing the real and $c$ the complex eigenvalues and ${N_{c}}$ the number of complex conjugated pairs. The operator $\mathbb{F}$ denotes the flip-operator, $\mathbb{F}(\sum_{i,j}c_{ij}{|{i,j}\rangle\!\rangle})=\sum_{i,j}c_{ij}^{*}{|{j,i}\rangle\!\rangle}$, which relates the spectral projectors of complex conjugated eigenvalue pairs. By taking the logarithm, we obtain a family of possible generators of $\mathcal{V}$ [54, 41],
$\hat{\mathcal{S}}_{\{\vec{x}\}}=\hat{\mathcal{S}}_{0}+i\frac{2\pi}{T}\sum_{c=1}^{N_{c}}x_{c}\left({|{r_{c}}\rangle\!\rangle}{\langle\!\langle{l_{c}}|}-\mathbb{F}{|{r_{c}}\rangle\!\rangle}{\langle\!\langle{l_{c}}|}\mathbb{F}\right),$ (3)
with $\hat{\mathcal{S}}_{0}$ stemming from the principal branch of the logarithm and a vector $\{\vec{x}\}=\{x_{1},x_{2},...,x_{N_{c}}\}\in\mathbb{Z}^{N_{c}}$ of integers labeling the possible branches of the complex logarithm. If all real eigenvalues $\lambda_{r}$ of $\hat{\mathcal{V}}$ are positive, all of these generators $\mathcal{S}_{\{\vec{x}\}}$ preserve hermiticity by construction and hence fulfil condition (i). At first glance, to test condition (ii), we have to inspect all branches, i.e., a countably infinite number of combinations of $N_{c}$ integers. Nevertheless, by use of methods from integer programming the problem can be reduced to a smaller set of possible integers [53], or alternatively one can apply machine-learning algorithms [57].
However it was shown [53] that deciding Markovianity is generally an NP-hard problem. This is especially detrimental in the case of interacting many-body systems where already the underlying Hilbert space $\mathcal{H}$ grows exponentially with system size $L$. Consider for example a spin chain of length $L$. Then, $\mathrm{dim}(\mathcal{H})=2^{L}$, so the maximum number of complex pairs (since $\lambda_{r}=1$ always exists [58]) is given by $N_{c}\leq 2^{2L-1}-1$. If we would just check the two closest branches around the principle branch, i.e. $x_{c}\in\{-1,0,1\}$, we already find super exponential scaling $3^{2^{2L-1}-1}$ of the number of branches for which condition (ii) has to be checked.
In the case that no branch gives rise to a valid Lindbladian, one can measure the distance to Markovianity and find the branch that is closest to a Markovian evolution. A possible measure proposed by Wolf et al. [54] that we use is based on adding a noise term $\chi\mathcal{Z}$ of strength $\chi$ to the generator $\mathcal{S}_{\{\vec{x}\}}$. The operator $\mathcal{Z}$ is the generator of the depolarizing channel $\exp({T\chi\mathcal{Z}})[\varrho]=e^{-\chi T}\varrho+(1-e^{-\chi T})\mathbf{1}/d$. The distance from Markovianity is defined as [54]
$\mu=\min_{\{\vec{x}\}\in\mathbb{Z}^{N_{c}}}\min\ \Bigl{\{}\chi\geq 0\big{|}\mathcal{S}_{\{\vec{x}\}}+\chi\mathcal{Z}\text{ is CCP}\Bigr{\}},$ (4)
the minimal strength needed, such that the generator $\mathcal{S}_{\{\vec{x}\}}+\chi\mathcal{Z}$ is Lindbladian.
Systems with access to full time-dependant map.—In the following we describe and test a simple strategy that allows to address (and often solve) this problem in cases, where additionally to $\mathcal{V}$ at time $T$, the full dynamical map $\mathcal{V}(t)$ is known also for times $t\in[0,T]$. To this end, we make connection to Floquet theory by continuing $\mathcal{V}(t)$ periodically for $t>T$, e.g. $\mathcal{V}(t)=\mathcal{V}(t-T)\mathcal{V}(T)$ for $T<t\leq 2T$ and so on. Then, the time evolution can effectively understood as generated from a time-periodic generator $\mathcal{G}(t)=\mathcal{G}(t+T)$ (not necessarily of Lindblad form) since
$\partial_{t}\varrho(t)=[\partial_{t}\mathcal{V}(t)]\mathcal{V}(t)^{-1}\varrho(t)=\mathcal{G}(t)\varrho(t),$ (5)
where we assume that $\mathcal{V}(t)$ is differentiable and invertible.
Let us briefly discuss Floquet theory for isolated systems. For systems with time-periodic Hamiltonian $H(t)=H(t+T)$, the fundamental solutions can be written in terms of Floquet states ${|{\psi_{\alpha}(t)}\rangle}=\exp(-i\varepsilon_{\alpha}t){|{u_{\alpha}(t)}\rangle}$ with time-periodic Floquet modes ${|{u_{\alpha}(t)}\rangle}={|{u_{\alpha}(t+T)}\rangle}$ and quasienergies $\varepsilon_{\alpha}$ [20]. Stroboscopically, at $t=0,T,2T,\dots$ the time evolution operator $U(T)=\exp(-iH_{\mathrm{F}}T)$ is described by the effective time-independent Floquet Hamiltonian $H_{\mathrm{F}}=\sum_{\alpha}\varepsilon_{\alpha}{|{u_{\alpha}(0)}\rangle}{\langle{u_{\alpha}(0)}|}$. Note that the quasienergies $\varepsilon_{\alpha}$ are not uniquely defined and can be redefined under the gauge transformation $\varepsilon_{\alpha}\rightarrow\varepsilon_{\alpha}+m\omega$ and ${|{u_{\alpha}(t)}\rangle}\rightarrow e^{im\omega t}{|{u_{\alpha}(t)}\rangle}$, where $\omega=2\pi/T$. Note that picking a specific gauge is equivalent to fixing the branch of the complex logarithm when solving for $H_{\mathrm{F}}=i\log[U(T)]/T$.
We now apply Floquet’s theorem to Eq. (5). The Floquet states take the form ${|{\varrho_{\mu}(t)}\rangle\!\rangle}=e^{-i\Omega_{\mu}t}{|{\Phi_{\mu}(t)}\rangle\!\rangle}$, where $\Omega_{\mu}$ are complex in general and the Floquet modes ${|{\Phi_{\mu}(t)}\rangle\!\rangle}$ are periodic in time [42, 58, 20]. Note that also here, redefining $\Omega_{\mu}\rightarrow\Omega_{\mu}+m\omega$ and ${|{\Phi_{\mu}(t)}\rangle\!\rangle}\rightarrow e^{im\omega t}{|{\Phi_{\mu}(t)}\rangle\!\rangle}$, with $m\in\mathbb{Z}$, leaves ${|{\varrho_{\mu}(t)}\rangle\!\rangle}$ invariant. In case of a purely coherent evolution, $\mathcal{G}(t)=-i[H(t),\cdot]$, the $\Omega_{\mu}$ correspond to quasienergy differences. In this paper, even in the case of dissipative dynamics, we still refer to the $\Omega_{\mu}$ as quasienergy differences, which are complex in general. For a given time $t$, we can decompose the dynamical map $\hat{\mathcal{V}}(t)$ as $\hat{\mathcal{V}}(t)=\sum_{\mu}e^{-i\Omega_{\mu}t}{|{\Phi_{\mu}(t)}\rangle\!\rangle}{\langle\!\langle{\tilde{\Phi}_{\mu}(0)}|}$, with the left Floquet modes ${\langle\!\langle{\tilde{\Phi}_{\mu}(t)}|}$ being different from the right ones for dissipative systems with a non-hermitian generator [42]. In this way, we can immediately identify the effective generator as [42] $\hat{\mathcal{S}}=-\sum_{\mu}i\Omega_{\mu}{|{\Phi_{\mu}(T)}\rangle\!\rangle}{\langle\!\langle{\tilde{\Phi}_{\mu}(0)}|},$ with the freedom of choosing the quasienergy differences $\Omega_{\mu}$ representing the freedom of choosing a branch of the logarithm in Eq. (3). Henceforth, we write ${|{\Phi_{\mu}(T)}\rangle\!\rangle}={|{\Phi_{\mu}(0)}\rangle\!\rangle}\equiv{|{\Phi_{\mu}}\rangle\!\rangle}$ and ${\langle\!\langle{\tilde{\Phi}_{\mu}(0)}|}\equiv{\langle\!\langle{\tilde{\Phi}_{\mu}}|}$.
Figure 1: (a)-(f) Distance from Markovianity $\mu$ (which here equals to distance from von-Neumann form $\mathcal{N}$) of the effective generator of the one-cycle evolution superoperator as a function of driving strength $E$ and frequency $\omega$, for no dissipation, $\gamma=0$, coupling strength $w=0.1$, and chain lengths of (a),(b),(d),(e) $L=2$ and (c),(f) $L=3$. To lift degeneracies in the spectrum, we set $\Delta_{1}=1.01,\Delta_{2}=1.00,\Delta_{3}=0.98$.
Let us now, for illustrative purposes regard the evolution with a time-independent generator $\partial_{t}\varrho(t)=\mathcal{L}\varrho(t)$, such that $\mathcal{V}(T)=e^{\mathcal{L}T}$, as an effective time-periodic problem with arbitrary period $T$. If we choose a large $T$ such that $\omega=2\pi/T$ is smaller than the imaginary part of some of the eigenvalues of $\mathcal{L}$, we observe that the generator $\mathcal{S}_{0}$ that we obtain from the principle branch of $\log(\mathcal{V}(T))/T$ is not identical to $\mathcal{L}$ but there occurs a “winding” of the eigenvalues of the Floquet generator $\mathcal{S}$ into the first Floquet Brillouin zone, $\mathrm{Re}\Omega_{\mu}\in[-\omega/2,\omega/2)$. This winding turns out to make the previously outlined procedure of searching around the principle branch inefficient, even for a fully coherent time evolution with no dissipation: To demonstrate this, we consider a circularly driven spin chain of length $L$
$H(t)=\sum_{\ell=1}^{L}\frac{\Delta_{\ell}}{2}\sigma_{z}^{\ell}+\sum_{\ell=1}^{L-1}w\sigma_{x}^{\ell}\sigma_{x}^{\ell+1}+H_{\text{drive}}(t)$ (6)
$H_{\text{drive}}(t)=\sum_{\ell=1}^{L}E\bigl{[}\cos(\omega t)\sigma_{x}^{\ell}-\sin(\omega t)\sigma_{y}^{\ell}\bigr{]},$ (7)
where $\Delta_{\ell}$ denotes the level splitting for the $\ell$-th spin and $w$ is the coupling strength between neighbouring spins. The driving is parameterized by the driving strength $E$ and -frequency $\omega$. As we show in Fig. 1(a) and (c), even if we calculate the dynamical map $\mathcal{V}(T)$ for an evolution with time-dependant von-Neumann generator $\mathcal{G}(t)=-i[H(t),\cdot]$, the generator $\mathcal{S}_{0}$ of the principle branch is not necessarily of von-Neumann form $\mathcal{N}$ (and may even be not of Lindblad form, cf. blue areas). This is in stark contrast to our knowledge of the existence of the Floquet Hamiltonian $H_{\mathrm{F}}$, i.e. $\mathcal{S}=-i[H_{\mathrm{F}},\cdot]$ is a valid effective Lindblad generator. Only after also searching around the first two neighbouring branches, i.e. $x_{c}\in\{-1,0,1\}$, one finds again a valid von-Neumann generator, as we show in Fig. 1(b). However, now, for every point in the phase diagram, we have a maximal number of $3^{2^{2L-1}-1}$ branches for which condition (ii) has to be checked. In the case of a spin chain of length $L=2$ this can still be computed in a reasonable time. However, due to the super exponential scaling, for $L=3$ ($L=4$) we would already have a maximal number of $\approx 6\cdot 10^{14}$ ($\approx 3\cdot 10^{60}$) branches to consider, simply to avoid the winding problem (cf. yellow line in Fig. 2(a)). Thus, in order to avoid this problem, we will now propose a different approach, which is based on a strategy for unwinding the quasienergy spectrum.
Spectral unwinding.—For undriven systems it is straight forward to define an “unwinding” procedure that undoes this winding into the first Floquet Brioullin zone: Since the Floquet modes ${|{\Phi_{c}(t)}\rangle\!\rangle}$ with complex quasienergy differences $\Omega_{c}$ are time-periodic, we can expand them in a Fourier series ${|{\Phi_{c}(t)}\rangle\!\rangle}=\sum_{n\in\mathbb{Z}}e^{i\omega nt}{|{\Phi_{c}^{(n)}}\rangle\!\rangle},$ and compute the trace norm for all of their Fourier components $||{|{\Phi_{c}^{(n)}}\rangle\!\rangle}||=\operatorname{Tr}\sqrt{\Phi_{c}^{\dagger(n)}\Phi_{c}^{(n)}}$. For every mode $c$ we determine $x_{c}^{(\text{max})}=\underset{x\in\mathbb{Z}}{\arg\max}\{||{|{\Phi_{c}^{(x)}}\rangle\!\rangle}||\}.$ Then, we shift the quasienergies $\Omega_{c}$ by this number of quanta $x_{c}^{\text{(max)}}$ which gives the ‘unwound’ Floquet generator $\hat{\mathcal{S}}_{\text{unf}}=\hat{\mathcal{S}}_{0}+i\frac{2\pi}{T}\sum_{c=1}^{N_{c}}x_{c}^{(\text{max})}\bigl{(}{|{\Phi_{c}}\rangle\!\rangle}{\langle\!\langle{\tilde{\Phi}_{c}}|}-\mathbb{F}{|{\Phi_{c}}\rangle\!\rangle}{\langle\!\langle{\tilde{\Phi}_{c}}|}\mathbb{F}\bigr{)}.$ In the undriven system, this gauge transformation removes the micromotion and transforms the Floquet modes into the static eigenmodes of the time-independent generator $\mathcal{G}$. In the following, we show that this procedure also significantly improves the search for an effective generator in the case of a time-dependent generator $\mathcal{G}(t)$. In this regard, we go back to our spin chain model. In Fig. 1(d) and (f) we show the phase diagrams we obtain for the generator of the principle branch $\mathcal{S}_{0}$ and the unwound generator $\mathcal{S}_{\text{unf}}$ for chain lengths $L=2$ and $L=3$, respectively. Here, between $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unf}}$, we always choose the generator with the smallest distance from Markovianity and plot its distance $\mu$. In the Supplemental Material (SM) [59], we plot the phase diagrams of $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unf}}$ seperately to show how they complement each other. By using the procedure above, in comparison to Fig. 1(a) and (c), where we have only considered $\mathcal{S}_{0}$, we find that specifically in the high-frequency regime we obtain a lot more points $(E,\omega)$ give generators of von-Neumann form. (In the SM [59], we introduce a measure to determine the distance from the von-Neumann form and show that both measures almost coincide for our problem.)
However, as we observe in Fig. 1(d),(f) for low frequencies, this method of unwinding fails. This is expected, since in this regime the peaks in the Fourier spectra can become more broadly distributed over several Fourier modes and indistinctly peaked. To overcome this issue, one has to include additional frequencies for the search in this regime. In the following, we outline how to find well-chosen candidates for effective generators to test for Markovianity, other than $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unf}}$. To this end, we set $x_{c}^{(\text{max,0})}=x_{c}^{(\text{max})}$ and chose a number $\eta\in[0,1]$, which determines at which ratio between the amplitudes of the neighbouring peaks at $x_{c}^{(\mathrm{max},i>0)}$ and the main peak at $x_{c}^{(\text{max,0})}$ we want to include them into the generation of branch combinations. The family of generators we consider therefore is $\hat{\mathcal{S}}_{\{\vec{\xi}\}}$ with the branches according to Eq. (3) but here, the integers $\xi_{c}$ are not any arbitrary integers, but the frequencies of the main and the included neighbouring peaks in the Fourier spectra of every Floquet mode
$\{\vec{\xi}\}=\Bigl{\{}(\xi_{i})\in\mathbb{Z}^{N_{c}}\ \big{|}$ $\xi_{c}\in\{x_{c}^{(\text{max,0})},...,x_{c}^{(\text{max,$z_{c}$})}\}\Bigr{\}},$ (8)
where for every pair of complex modes $c$, we determine the number $z_{c}\in\mathbb{N}$ which counts how many neighbouring peaks have amplitudes that are on the order of the amplitude of the main peak according to the criterion ${A^{(\text{max,i})}_{c}}/{A^{(\text{max,0})}_{c}}\geq\eta,$ with $A^{(\text{max},i)}_{c}=\operatorname{Tr}\sqrt{\Phi_{c}^{\dagger(x^{(\text{max,i})})}\Phi_{c}^{(x^{(\text{max,i})})}}$. We additionally introduce a cutoff $N_{b}$ for the maximal number of branches considered and only consider the peaks $i=1,...,N_{b}$. This gives us a way to systematically search around the unwound generator $\mathcal{S}_{\text{unf}}$, instead of searching around the generator of the principle branch $\mathcal{S}_{0}$, which can drastically reduce the required numerical effort. In Fig. 1(e) we apply this modified Markovianity test for a spin chain of length $L=2$ and the parameters $N_{b}=2$ and $\eta=0.7$. We observe that we almost fully recover valid von-Neumann generators. Note that for $L\geq 3$ this choice of $N_{b}$ and $\eta$ will not suffice to recover valid von-Neumann generators in every point of the phase diagram. Here, to minimise the complexity, for a given point one can search for the smallest $N_{b}$ and largest $\eta$ which yield a generator of von-Neumann form. In the SM [59], for $L=3$ we have set $N_{b}=1$ and varied $\eta$ to find an optimum along the vertical red line at $E=2.5$ in Fig. 1(e).
Figure 2: (a) Cardinality $|M|$ of the set $M$ of possible branch combinations for the different methods and lengths $L$. (b)-(e) Distance from Markovianity $\mu$ of the effective generator of the one-cycle evolution superoperator as a function of driving strength $E$ and frequency $\omega$, for weak dissipation $\gamma=0.01$, bath temperature $\mathcal{T}=1$ and chain lengths of (b),(d) $L=3$ and (c),(e) $L=4$. Other parameters $\Delta_{1}=1.4,\Delta_{2}=1.1,\Delta_{3}=0.7,\Delta_{4}=0.9,w=0.01$.
To compare the reduction in complexity of our method to the naive procedure of searching around the principle branch, in Fig. 2(a) we plot the cardinality $|M|$ of the set $M$ of all possible branch combinations for which we test for Markovianity. For the standard Markovianity test the size of $M$ is dependent on the included neighboring branches $x_{c}$. For the modified Markovianity test the size of $M$ depends on our choice of $N_{b}$ and $\eta$, which we choose such that we minimise the included branches but still get a proper von-Neumann generator. For the principle branch $\mathcal{S}_{0}$ we always have $|M|=1$ (blue line). The generators of the principle branch yields poor results due to the winding of quasienergy differences. Including the first two neighbouring, i.e. $x_{c}\in\{-1,0,1\}$, fixes this problem, but greatly increases the complexity as we see for the orange line. On the other hand, trying to revert the winding by only including $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unf}}$, we have $|M|=2$ (green line). Using the modified Markovianity test of Eq. (8), the minimal cardinality needed to obtain a fully von-Neumann phase diagram for $L=2$ and to make the points on the red line in Fig. 1(c) fully von-Neumann for $L=3$, is plotted by the red line in Fig. 2(a). For a spin chain of length $L=3$ we have $|M|\approx 10^{5}$ on average (see SM [59] for details) to make the red line fully von-Neumann. Searching around the principle branch requires $|M|\approx 10^{13}$. This shows that searching around the unwound generator dramatically decreases the complexity (as long as we are not at low frequencies). Nevertheless, even for the modified method, the complexity still grows fast with the dimensionality of the problem, which is why for bigger systems one might only be able to include the generator of the principle branch $\mathcal{S}_{0}$ and the unwound generator $\mathcal{S}_{\text{unf}}$ (or alternatively, combine our method with the existing machine learning method of Ref. [57]).
Dissipative case.—In the case of a coherent time evolution we can always compute a Floquet Hamiltonian $H_{\mathrm{F}}$ and then obtain the effective time-independent Floquet generator $\mathcal{S}=-i[H_{\mathrm{F}},\cdot]$, which always has von-Neumann form. Such a bypass, however, is not possible in the case of dissipative systems where we search for effective time-independent generators of Lindblad form, i.e. Floquet Lindbladians [41, 42], since here we have to compute $\log(\mathcal{V}(T))/T$ and test these generators for Markovianity [41]. If we do not change the Hamiltonian but add very weak dissipation, we can see exactly the same non-Markovian areas arising in the principle branch (due to the winding of the eigenvalues of $\log[\mathcal{V}(T)]/T$), cf. Fig. 2(b). Here, we add dissipation described by Lindblad jump operators for the quantum-optical master equation $L^{\ell}_{kq}=\sqrt{R^{\ell}_{kq}}{|{\psi_{k}}\rangle}{\langle{\psi_{q}}|},$ where, for simplicity, we neglect the impact of the driving term on the dissipator. Hence, ${|{\psi_{k}}\rangle},E_{k}$ are the many-body eigenstates and -energies of the undriven Hamiltonian for $E=0$. The corresponding jump rate for contact with a reservoir at temperature $\mathcal{T}$ at site $l$ reads $R^{\ell}_{kq}=2\pi\gamma^{2}|{\langle{\psi_{k}}|}\sigma_{x}^{\ell}{|{\psi_{q}}\rangle}|^{2}g(E_{k}-E_{q})$ with $g(E)=E/(e^{E/\mathcal{T}}-1)$, assuming an ohmic bath [60]. By summing over all jump operators we obtain the total time-periodic Lindbladian $\mathcal{L}(t)[\cdot]=-{i}\bigl{[}H(t),\cdot\bigr{]}+\sum_{\ell=1}^{L}\sum_{k,q}\bigl{(}L^{\ell}_{kq}\cdot L_{kq}^{\ell\dagger}-\frac{1}{2}\bigl{\{}L_{kq}^{\ell\dagger}L^{\ell}_{kq},\cdot\bigr{\}}\bigr{)}.$ Only considering the generator of the principle branch $\mathcal{S}_{0}$, for $L=3$ and $L=4$, we obtain the map shown in Fig. 2(b) and (c), respectively. Applying the Markovianity test to the unwound Floquet generator $\mathcal{S}_{\text{unf}}$ and $\mathcal{S}_{0}$ yields the phase diagrams in Fig. 2(d) and (e), where we observe that the phase diagrams contain more valid Floquet Lindbladians $\mathcal{L}_{\mathrm{F}}$ (white region) [41, 42] or have significantly lower distances from Markovianity $\mu$ in the areas where no $\mathcal{L}_{\mathrm{F}}$ exists. Note again that scanning through all the possible logarithm branches would mean to scan through possibly $\approx 10^{14}$ candidates for $L=3$ and $\approx 10^{60}$ candidates for $L=4$, if we merely wanted to include the first two neighbouring branches $x_{c}\in\{-1,0,1\}$. The unwound generator $\mathcal{S}_{\text{unf}}$ thus gives rise to a fast way of determining a particularly good candidate generator which we can test for Markovianity, when dealing with high-dimensional open quantum many-body systems. In the Supplemental Material we give an example to show that the unwound generator is a good candidate not only for Floquet systems but also in the general case of a time-dependent non-periodic map $\mathcal{V}(t)$. We show that for a convex combination $\mathcal{V}(t)=\lambda\mathcal{V}_{1}(t)+(1-\lambda)\mathcal{V}_{2}(t)$ [61], where $\mathcal{V}_{j}(t)=e^{t\mathcal{L}_{j}}$ and $\lambda\in[0,1]$, the unwound generator $\mathcal{S}_{\text{unf}}$ produces valid effective Lindbladians in areas where the generator of principle branch $\mathcal{S}_{0}$ does not.
Summary.—We have demonstrated that the problem of deciding Markovianity [54] (i.e. deciding whether a given quantum channel was generated by a Lindbladian generator) can often be solved for cases with access to the full dynamical map. To this end, we employ Floquet theory and perform a gauge transformation that minimizes the occurring micromotion. Our ideas also significantly simplify the calculation of effective Floquet-Lindbladians for complex many-body systems [41, 42, 57] and can be applied in the context of Liouvillian learning [27] for trotterized open quantum systems. Remarkably, already without dissipation, we have shown that the effective Floquet generator is generally not of von-Neumann form. This highlights the fact that while for the Hamiltonian case any branch of the quasienergy spectrum yields a sensible description, in the case of a Liouvillian one has to be much more careful, since only some branches provide a valid physical description.
Acknowledgements.